# GRAPES: Learning to Sample Graphs for Scalable Graph Neural Networks

### v2

@@ 5 Experiments
Our experiments aim at answering the following research question: given a fixed sampling budget and GNN architecture, what is the effect of training with an adaptive policy for layer-wise sampling, in comparison with the related work? While several works in the literature of sampling for GNNs have focused on classification benchmarks to demonstrate the performance of sampling algorithms, there are several confounding factors in their experimental setup that prevent proper understanding whether the perceived performance improvements are a result of the sampling method itself. Examples of such confounding factors found in related work include using different architectures, like the GCN [23] or GAT [38], different sizes for the hidden layers, number of layers, batch sizes, number of nodes sampled per layer, and the number of training epochs. These are factors independent of sampling algorithms that nonetheless affect the performance in the benchmarks. We present a detailed overview of differences in the experimental setup of the related work in Table 3 in the appendix.
To better understand the effect of sampling algorithms, we thus carry out experiments on a fixed GCN architecture [23] with two layers, a hidden size of 256, batch size of 256, sample size of 256 nodes per layer, and a fixed number of epochs per dataset (detailed in Appendix B). Under this setting, we compare GRAPES with the following baselines: a Random baseline that uses the same setup as GRAPES but with uniform inclusion probabilities; FastGCN [9], LADIES [51], GraphSAINT [47], GAS [13], and AS-GCN [18]. With this setup, we aim to control our experiments in a way such that variations in performance can only be attributed to the sampling method.
For all baselines (except Random) we rely on their publicly available implementations. For all methods, we optimize the learning rate using the performance on the validation set. GRAPES requires selecting two additional hyperparameters: the learning rate for the sampling policy and, for GRAPES-GFN, the reward scaling parameter $\alpha$. We tune these based on the classification performance on the validation set. These hyperparameters are specific to our sampling method and do not explicitly affect the learning capacity of the GCN classifier.
We refer the reader to Appendix B for more details on hyperparameter settings.
@@ 5.1 Datasets
Most methods in the literature of sampling in GNNs are evaluated on homophilous graphs (where similar nodes are likely to be connected by an edge) and multi-class classification, where nodes are classified into one of several classes (in contrast to multi-label). To further understand the effect of sampling, we carry out experiments with heterophilous graphs, where nodes that are connected differ in their features and labels, and multi-label datasets in which a node can be assigned one or more labels. As shown in the literature [50, 48], we argue that node classification with GCN is more challenging on these datasets compared to multi-class classification on homophilous graphs, where an adaptive policy could be beneficial. In particular, we run experiments with the following datasets for the node classification task:
  • Homophilous graphs: citation networks (Cora, Citeseer, Pubmed with the “full” split) [34, 43], Reddit [16], ogbn-arxiv and ogbn-products [17], and DBLP [48].
  • Heterophilous graphs: Flickr [47], Yelp [47], ogbn-proteins [17], BlogCat [48], and snap-patents [24] .
We included the label homophily ratio of all the datasets in Table 1 and 2. For the multi-label datasets, we followed the same framework in [48] to calculate their label homophily ratio. Statistics about the datasets can be found in Appendix C.
Table 1: F1-scores (%) for different sampling methods trained on homophilous graphs, for a batch size of 256, and sample size of 256 per layer. We report the mean and standard deviation over 10 runs. The best values among the sampling baselines (all except GAS) are in bold, and the second best are underlined. MC stands for multi-class and ML stands for multi-label classification.
@@
Table 2: F1-scores (%) for different sampling methods trained on heterophilous graphs for a batch size of 256, and a sample size of 256 per layer. We report the mean and standard deviation over 10 runs. The best values among the sampling baselines (all except GAS) are in bold, and the second best are underlined. MC stands for multi-class and ML stands for multi-label classification.
@@
GRAPES-RL (ours) 49.54 $\pm$ 0.67 26.26 $\pm$ 0.50 40.69 $\pm$ 0.55 11.78 $\pm$ 0.14 9.06 $\pm$ 0.68
GRAPES-GFN (ours) 49.29 $\pm$ 0.32 28.46 $\pm$ 0.69 44.57 $\pm$ 0.88 11.57 $\pm$ 0.18 9.22 $\pm$ 0.40
@@ 5.2 Results
Comparison with sampling methods We present the F1 scores of GCNs trained via sampling for GRAPES and the sampling-based baselines in Table 1 and 2. We observe that on the majority of heterophilous datasets, and all the heterophilous multi-label datasets, both variations of GRAPES achieve the highest F1-score. As we mentioned earlier, node classification on heterophilous graphs is a challenging task for GCNs, due to the diversity of the neighbor nodes’ labels. Our results suggest that GRAPES’ ability to adapt the sampling policy to particular features of the data allows GRAPES to learn the complex patterns across the heterophilous datasets. Table 1 shows that on most of the homophilous graphs, AS-GCN achieves the highest F1-score, indicating the importance of adaptive sampling. However, unlike GRAPES, this method is only adaptive to the sampling variance and fails to outperform GRAPES on the heterophilous graphs.
FastGCN and LADIES, which use a fixed policy, fail to compete with the other methods in our experimental setup. They assign a higher probability to nodes with higher degrees. This heuristic results in neglecting informative low-degree nodes. Although outperforming LADIES and FastGCN, GraphSAINT, which also uses a fixed sampling policy, fails to outperform the more competitive baselines.
We observe that on some datasets, Random sampling achieves a comparable F1-score as GRAPES and the baselines. This mainly happens on simple datasets, such as homophilous graphs, indicting that on these graphs a simple sampling approach is sufficient.
Comparison with a non-sampling scalable method GAS [13] is a non-sampling method that uses the historical embeddings of the 1-hop neighbors of the target nodes. We present classification results in Table 2, and we visualize memory usage in Fig. 2. While GAS has higher classification F1-scores for some datasets, GRAPES achieves comparable F1-scores and significantly outperforms GAS for Yelp, ogbn-proteins, and BlogCat, all heterophilous multi-label graphs. Once again, these results indicate the effectiveness of the adaptivity of GRAPES in sampling influential nodes in complex graphs. For the memory usage, we compare GAS with the GFlowNet version of GRAPES-32 and GRAPES-256, where $32$ and $256$ (the default value) indicate the sample size per layer. The memory usage results show that, even with a large sample size, GRAPES can use up to an order of magnitude less GPU memory than GAS, especially for large datasets. In large, densely connected graphs such as Reddit, the 1-hop neighborhood can be massive. Then, the difference in memory use for GRAPES, which only sees a small set of neighbors, and GAS, which uses all the neighbors, is significant. While GAS occasionally achieves higher F1-scores, it consistently demands more memory, indicating a potential compromise between accuracy and computational efficiency. In contrast, GRAPES strikes a balance, delivering comparable F1-score with more modest memory footprints.
@@
GRAPES learns strong preferences over nodes. The ability to selectively choose influential nodes is a crucial property of GRAPES. Figure 4 shows the mean and standard deviation of base 2 entropy for the node preference probabilities for the two layers of $\operatorname{GCN}_{\operatorname{S}}$ for ogbn-products and DBLP for GRAPES-GFN. The probabilities show preference towards particular nodes with a Bernoulli distribution. A well-trained model must have a high preference (probability close to 1) for some nodes and a low preference (probability close to 0) for the rest. Therefore, we would like a low average entropy with a high standard deviation. As the figure shows, the mean entropy in both layers decreases from almost 1 and converges to a value above 0 while the standard deviation increases. This indicates that for ogbn-products, the sampling policy initially assigned a probability near $0.5$, indicating little preference. However, after several training epochs, GRAPES starts preferring some nodes, resulting in lower mean entropy. We observe similar behavior for most datasets (DBLP, BlogCat, Yelp, ogbn-arxiv, and ogbn-proteins). However, we observe that the sampling policy exhibits no preferences among the nodes for the other datasets. For more details about the other datasets, see Appendix D.
@@
6 Discussion and Conclusion
We propose GRAPES, an adaptive graph sampling method based on reinforcement learning and GFlowNet, facilitating the scalability of training GNNs on massive graphs. GRAPES learns the importance of the nodes by adapting to the classification loss. We show how a sampling policy can build a subgraph of influential nodes by learning node preferences that adapt to node features, GNN architecture, classification task, and graph topology. Our experiments demonstrate that GRAPES effectively selects nodes from large-scale graphs and achieves consistent performance over state-of-the-art sampling methods, especially on complex graph structures such as heterophilous graphs. Compared to the other sampling methods, GRAPES can maintain high classification accuracy even with lower sampling rates, indicating GRAPES’ ability to scale to larger graphs by sampling a small but influential set of nodes. GRAPES achieves comparable performance to GAS, while using up to an order of magnitude less memory.
Lack of Uniform Evaluation Protocol. Existing methods in the literature report performance on graph sampling under settings with different GCN architectures, regularization techniques, feature normalization strategies, and data splits, among other differences. These differences made it challenging to determine the benefits of each sampling method. This motivated us to implement a unified protocol across all methods, where we keep the architecture fixed. We encourage future work to consider a similar methodology for a fair evaluation, or an experimental review study, as is common in other areas of machine learning research on graphs [36, 32].
@@ 6 Discussion and Conclusion
Known Limitations. In our experiments, we only focused on the problem of node classification. However, GRAPES is not tied to a particular downstream task. GRAPES assumes access to a tractable reward function [7]. Therefore, GRAPES will be applicable for other graph-related tasks, like link prediction and unsupervised representation learning. Another limitation is that we only evaluate GRAPES on a GCN architecture with a fixed number of layers. Additionally, there is room for a more thorough investigation into the properties of the subgraphs that GRAPES samples. The experiments at the end of Section 5.2 are a first step towards this. We leave these directions for future work.
Acknowledgments
@@
This project was partially funded by Elsevier’s Discovery Lab and Huawei DREAMS Lab. Emile van Krieken was funded by ELIAI (The Edinburgh Laboratory for Integrated Artificial Intelligence), EPSRC (grant no. EP/W002876/1). We thank Michael Cochez and Ruud van Bakel for insightful discussions. We also thank the Das6 cluster for computing support [5].
@@ Appendix A Off-Policy Sampling Setup
In this Appendix, we discuss the technical and mathematical challenges around our setup that resulted in our off-policy learning setup. In each layer $l$ of the GFlowNet, we aim to sample exactly $k$ out of $n$ nodes. An initially natural setup would be to use the distribution over $k$-subsets of $\mathcal{N}(K^{(l-1)})$ [3]. Using Bayes theorem,
@@
When conditioned on $k$, $q$ assigns 0 probability to sets of nodes $\mathcal{V}^{(l)}$ that do not sample exactly $k$ new nodes (that is, when $|\mathcal{V}^{(l)}|\neq k$). However, this requires renormalizing the distribution, which is the function of the denominator term on the right-hand side. Note that this sum is over an exponential number of elements, namely $2^{|\mathcal{N}(K^{(l-1)})|}$, and naive computation is clearly intractable. SIMPLE [3] provides an optimized dynamic programming algorithm for computing this normalization constant. However, it scales polynomially in $|\mathcal{N}(K^{(l-1)})|$ and $k$, and in our experiments, computing the normalizer is a bottleneck already for mid-sized graphs like Reddit.
Therefore, we decided to circumvent having to compute $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)},k)$ by sampling using the Gumbel-Top-k trick (Equation 4) to ensure we always add exactly $k$ nodes. However, we are now in an off-policy setting: The samples using Equation 4 are distributed by $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)},k)$, not by $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})$, and so we sample from a different distribution than the one we use to compute the loss. Previous work [29] showed that the Trajectory Balance loss is amenable to off-policy training without importance sampling and weighting without introducing high variance. This is important since importance weighting would require us to weight by $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})/q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)},k)$, reintroducing the need to compute $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)},k)$.

### v3

@@ 6 Experiments
Our experiments aim at answering the following research question: given a fixed sampling budget and GNN architecture, what is the effect of training with an adaptive policy for layer-wise sampling, in comparison with the related work? While several works in the literature of sampling for GNNs have focused on classification benchmarks to demonstrate the performance of sampling algorithms, several confounding factors in their experimental setup prevent a proper understanding of whether the perceived performance improvements result from the sampling method itself. Examples of such confounding factors found in related work include using different architectures, like the GCN (Kipf & Welling, 2016) or GAT (Velickovic et al., 2017), different sizes for the hidden layers, number of layers, batch sizes, number of nodes sampled per layer, and the number of training epochs. These are factors independent of sampling algorithms that nonetheless affect the performance in the benchmarks. We present a detailed overview of differences in the experimental setup of the related work in Table 4 in the appendix.
To better understand the effect of sampling algorithms, we thus carry out experiments on a fixed GCN architecture (Kipf & Welling, 2016) with two layers, a hidden size of 256, batch size of 256, sample size of 256 nodes per layer, and a fixed number of epochs per dataset (detailed in Appendix B). Under this setting, we compare GRAPES with the following baselines: a Random baseline that uses the same setup as GRAPES but with uniform inclusion probabilities; FastGCN (Chen et al., 2018b), LADIES (Zou et al., 2019), GraphSAINT (Zeng et al., 2019), GAS (Fey et al., 2021), AS-GCN (Huang et al., 2018), and PASS Yoon et al. (2021). With this setup, we aim to control our experiments in a way such that variations in performance can only be attributed to the sampling method.
For all baselines (except Random), we rely on their publicly available implementations. For all methods, we optimize the learning rate using the performance on the validation set. GRAPES requires selecting two additional hyperparameters: the learning rate for the sampling policy and, for GRAPES-GFN, the reward scaling parameter $\alpha$. We tune these based on the classification performance on the validation set. These hyperparameters are specific to our sampling method and do not explicitly affect the learning capacity of the GCN classifier. We refer the reader to Appendix B for more details on hyperparameter settings.
@@ 6.1 Datasets
Most methods in the literature of sampling in GNNs are evaluated on homophilous graphs (where the connected nodes are likely to have similar labels) and multi-class classification, where nodes are classified into one of several classes (in contrast to multi-label). To further understand the effect of sampling, we carry out experiments with heterophilous graphs, where nodes that are connected differ in their features and labels, and multi-label datasets in which a node can be assigned one or more labels. As shown in the literature (Zhu et al., 2020; Zhao et al., 2023), we argue that node classification with GCN is more challenging on these datasets compared to multi-class classification on homophilous graphs, where an adaptive policy could be beneficial. In particular, we run experiments with the following datasets for the node classification task:
  • Homophilous graphs: citation networks (Cora, Citeseer, Pubmed with the “full” split) (Sen et al., 2008; Yang et al., 2016), Reddit (Hamilton et al., 2017), ogbn-arxiv and ogbn-products (Hu et al., 2020), and DBLP (Zhao et al., 2023).
  • Heterophilous graphs: Flickr (Zeng et al., 2019), Yelp (Zeng et al., 2019), ogbn-proteins (Hu et al., 2020), BlogCat (Zhao et al., 2023), and snap-patents (Leskovec & Krevl, 2014).
We included the edge homophily ratio (Zhu et al., 2020) of all the datasets in Table 1 and 2. For the multi-label datasets, we followed the same framework in (Zhao et al., 2023) to calculate their label homophily ratio. Statistics about the datasets can be found in Appendix C.
Table 1: F1-scores (%) for different sampling methods trained on homophilous graphs, for a batch size of 256, and sample size of 256 per layer. We report the mean and standard deviation over 10 runs. The best values among the sampling baselines (all except GAS) are in bold, and the second best are underlined. MC stands for multi-class and ML stands for multi-label classification. OOM indicates out-of-memory.
@@
Rank 1 2 3 1 3 2 2
Table 2: F1-scores (%) for different sampling methods trained on heterophilous graphs for a batch size of 256, and a sample size of 256 per layer. We report the mean and standard deviation over 10 runs. The best values among the sampling baselines (all except GAS) are in bold, and the second best are underlined. MC stands for multi-class and ML stands for multi-label classification. OOM indicates out of memory.
@@
GRAPES-RL (ours) 49.54 $\pm$ 0.67 29.16 $\pm$ 0.54 40.69 $\pm$ 0.55 11.78 $\pm$ 0.14 9.06 $\pm$ 0.68
GRAPES-GFN (ours) 49.29 $\pm$ 0.32 29.58 $\pm$ 0.24 44.57 $\pm$ 0.88 11.57 $\pm$ 0.18 9.22 $\pm$ 0.40
Rank 1 4 1 1 1
@@ 6.2 Results
Comparison with sampling methods We present the F1 scores of GCNs trained via sampling for GRAPES and the sampling-based baselines in Table 1 and 2. We observe that on the majority of heterophilous datasets, and all the heterophilous multi-label datasets, both variations of GRAPES achieve the highest F1-score. As we mentioned earlier, node classification on heterophilous graphs is a challenging task for GCNs, due to the diversity of the neighbor nodes’ labels. Our results suggest that GRAPES’ ability to adapt the sampling policy to particular features of the data allows GRAPES to learn the complex patterns across the heterophilous datasets. Table 1 shows that on most homophilous graphs, AS-GCN achieves the highest F1-score, indicating the importance of adaptive sampling. However, unlike GRAPES, this method is only adaptive to the sampling variance and fails to outperform GRAPES on the heterophilous graphs. Moreover, our results show that PASS, another adaptive method, is unable to compete with GRAPES and results in an out-of-memory error on several large datasets. This is due to its node-wise sampling feature, which results in a disconnected sampled graph.
FastGCN and LADIES, which use a fixed policy, fail to compete with the other methods in our experimental setup. They assign a higher probability to nodes with higher degrees. This heuristic results in neglecting informative low-degree nodes. Although outperforming LADIES and FastGCN, GraphSAINT, which also uses a fixed sampling policy, fails to outperform the more competitive baselines. We observe that on some datasets, Random sampling achieves a comparable F1-score to GRAPES and the baselines. This mainly happens on simple datasets, such as homophilous graphs, indicating that on these graphs, a simple sampling approach is sufficient.
Comparison with a non-sampling scalable method GAS (Fey et al., 2021) is a non-sampling method that uses the historical embeddings of the 1-hop neighbors of the target nodes. We present classification results in Table 2, and we visualize memory usage in Fig. 3. While GAS has higher classification F1-scores for some datasets, GRAPES achieves comparable F1-scores and significantly outperforms GAS for Yelp, ogbn-proteins, and BlogCat, all heterophilous multi-label graphs. Once again, these results indicate the effectiveness of the adaptivity of GRAPES in sampling influential nodes in complex graphs. For the memory usage, we compare GAS with AS-GCN and the GFlowNet version of GRAPES-32 and GRAPES-256, where $32$ and $256$ (the default value) indicate the sample size per layer.
@@
The memory usage results show that, even with a large sample size, GRAPES can use up to an order of magnitude less GPU memory than GAS, especially for large datasets. In large, densely connected graphs such as Reddit, the 1-hop neighborhood can be massive. Then, the difference in memory use for GRAPES, which only sees a small set of neighbors, and GAS, which uses all the neighbors, is significant. While GAS occasionally achieves higher F1-scores, it consistently demands more memory, indicating a potential compromise between accuracy and computational efficiency. In contrast, GRAPES strikes a balance, delivering comparable F1-score with more modest memory footprints. In most cases, GRAPES also uses less memory than AS-GCN, another sampling-based baseline. In other cases, GRAPES uses more memory than AS-GCN, which we attribute to the fact that there is an additional GNN in GRAPES for learning the sampling policy. Some variations are due to implementation details [^4] We used PyTorch for our implementation, whereas AS-GCN is implemented in Tensorflow, which manages GPU memory differently., though in general we observe a significant advantage of layer-wise sampling methods over historical embeddings.
@@
GRAPES learns strong preferences over nodes. The ability to selectively choose influential nodes is a crucial property of GRAPES. Figure 5 shows the mean and standard deviation of base 2 entropy for the node preference probabilities for the two layers of $\operatorname{GCN}_{\operatorname{S}}$ for ogbn-products and DBLP for GRAPES-GFN. The probabilities show preference towards particular nodes with a Bernoulli distribution. A well-trained model must have a high preference (probability close to 1) for some nodes and a low preference (probability close to 0) for the rest. Therefore, we would like a low average entropy with a high standard deviation. As the figure shows, the mean entropy in both layers decreases from almost 1 and converges to a value above 0, while the standard deviation increases. This indicates that for ogbn-products, the sampling policy initially assigned a probability near $0.5$, indicating little preference. However, after several training epochs, GRAPES starts preferring some nodes, resulting in lower mean entropy. We observe similar behavior for most datasets (DBLP, BlogCat, Yelp, ogbn-arxiv, and ogbn-proteins). However, we observe that the sampling policy exhibits no preferences among the nodes for the other datasets. For more details about the other datasets, see Appendix D.
@@
7 Known Limitations and Future Work
Our experimental results show that GRAPES outperforms the baselines on some, but not all, heterophilous datasets. We provide a theoretical analysis that shows the effectiveness of adaptive sampling on certain types of heterophilous graphs. Further analysis is required to understand which graph characteristics contribute to the performance gain observed with adaptive sampling methods such as GRAPES.
@@ 7 Known Limitations and Future Work
Moreover, in our experiments, we focused only on the problem of node classification. However, GRAPES is not tied to a particular downstream task. GRAPES assumes access to a tractable reward function (Bengio et al., 2021b). Therefore, GRAPES will be applicable for other graph-related tasks, like link prediction and unsupervised representation learning.
Additionally, there is room for a more thorough investigation of the properties of the subgraphs that GRAPES samples. The experiments at the end of Section 6.2 and in Appendix D are a first step towards this. We leave these directions for future work.
Another direction for future work is applying GRAPES to architectures designed to address heterophily, such as the ones proposed by Zhu et al. (2020); Abu-El-Haija et al. (2019). By leveraging GRAPES, these architectures could potentially overcome their scalability issues.
8 Discussion, Broader Impacts, and Conclusion
We propose GRAPES, an adaptive graph sampling method based on reinforcement learning and GFlowNet, facilitating the scalability of training GNNs on massive graphs. GRAPES samples a subgraph of influential nodes by learning node preferences that adapt to classification loss, which depends on node features, GNN architecture, classification task, and graph topology. Our experiments demonstrate that GRAPES effectively selects nodes from large-scale graphs and achieves state-of-the-art performance in multi-label classification tasks on heterophilous graphs, while performing competitively for multi-class classification on homophilous graphs. Compared to the other sampling methods, GRAPES can maintain high classification accuracy even with lower sample sizes, indicating GRAPES’ ability to scale to larger graphs by sampling a small but influential set of nodes. GRAPES achieves comparable performance to GAS, while using up to an order of magnitude less memory.
Focus on Heterophily in Sampling. Previous works have overlooked the impact of sampling on heterophilous and multi-label graphs. To our knowledge, this is the first work to compare sampling effects on both homophilous and heterophilous graphs, where adaptive sampling is especially crucial due to the diversity among neighbors, making node selection vital for GNN accuracy.
Lack of Uniform Evaluation Protocol. Existing methods in the literature report performance on graph sampling under settings with different GCN architectures, regularization techniques, feature normalization strategies, and data splits, among other differences. These differences made it challenging to determine the benefits of each sampling method. This motivated us to implement a unified protocol across all methods, where we keep the architecture fixed. We encourage future work to consider a similar methodology for a fair evaluation, or an experimental review study, as is common in other areas of machine learning research on graphs (Shchur et al., 2018; Ruffinelli et al., 2019).
Broader Impacts. Our paper focuses on fundamental research to understand and advance sampling in large-scale graphs. There may be multiple potential societal consequences of our work, but there are none that can be easily predicted and specifically highlighted.
Acknowledgements
@@
Taraneh Younesian was funded by Huawei DREAMS Lab. All content represents the opinion of the authors, which is not necessarily shared nor endorsed by their respective employers and/or sponsors in Huawei DREAMS Lab. Daniel Daza was partially funded by Elsevier’s Discovery Lab. Emile van Krieken was funded by ELIAI (The Edinburgh Laboratory for Integrated Artificial Intelligence), EPSRC (grant no. EP/W002876/1). We thank Michael Cochez and Ruud van Bakel for insightful discussions.
@@ Appendix A Off-Policy Sampling Setup
In this Appendix, we discuss the technical and mathematical challenges around our setup that resulted in our off-policy learning setup. In each layer $l$ of the GFlowNet, we aim to sample exactly $k$ out of $n$ nodes. An initially natural setup would be to use the distribution over $k$-subsets of $\mathcal{N}(K^{(l-1)})$ (Ahmed et al., 2023). Using Bayes theorem,
@@
When conditioned on $k$, $q$ assigns 0 probability to sets of nodes $\mathcal{V}^{(l)}$ that do not sample exactly $k$ new nodes (that is, when $|\mathcal{V}^{(l)}|\neq k$). However, this requires renormalizing the distribution, which is the function of the denominator term on the right-hand side. Note that this sum is over an exponential number of elements, namely $2^{|\mathcal{N}(K^{(l-1)})|}$, and naive computation is clearly intractable. SIMPLE (Ahmed et al., 2023) provides an optimized dynamic programming algorithm for computing this normalization constant. However, it scales polynomially in $|\mathcal{N}(K^{(l-1)})|$ and $k$, and in our experiments, computing the normalizer is a bottleneck already for mid-sized graphs like Reddit.
Therefore, we decided to circumvent having to compute $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)},k)$ by sampling using the Gumbel-Top-k trick (Equation 4) to ensure we always add exactly $k$ nodes. However, we are now in an off-policy setting: The samples using Equation 4 are distributed by $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)},k)$, not by $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})$, and so we sample from a different distribution than the one we use to compute the loss. Previous work (Malkin et al., 2022b) showed that the Trajectory Balance loss is amenable to off-policy training without importance sampling and weighting without introducing high variance. This is important since importance weighting would require us to weight by $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)})/q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)},k)$, reintroducing the need to compute $q(\mathcal{V}^{(l)}|\mathcal{V}^{(0)},\ldots,\mathcal{V}^{(l-1)},k)$.